# Stability theory of flat band solitons in nonlinear wave systems

### v1

@@ Abstract
It is therefore particularly relevant to study the types of stable or long-lived localized excitations that can be supported in nonlinear wave models, whose underlying linear spectrum has a flat band. A prototypical example arises in the discrete nonlinear Schrödinger (DNLS) type wave systems arising in optics, superfluids and materials (see e.g. [Lederer2008_PhysRep_DiscreteSolitons, Kevrekidis2009_DNLS, malomed2013spontaneous, AblowitzCole2022_PhysicaD_Review, granularBook]), on the Kagomé lattice [Law]. Numerical studies reveal compactly supported nonlinear bound states from which one observes symmetry-breaking bifurcations  [vicencio2013discrete]. While there is substantial analytical theory of solitary wave (“soliton”) stability, most information about nonlinear dynamics of coherent structures in systems with linear flat band spectrum has been obtained by direct simulation. A general, unifying theory of the dynamical stability of the nonlinear states of such systems, is of potential broad interest and our goal in this Letter is to present such a theory.
@@
The band structure contains a flat band if for some $j_{0}\in\{1,\dots,N\}$, $E_{j_{0}}(k)=E_{F}$ is a constant for all $k\in\mathcal{B}$. In this paper, we focus on the setting where $\Delta_{\mathbb{L}}$ has a single flat band at the upper limit, $E_{F}$, of its spectrum. The general properties of linear operators with flat bands have been studied extensively. For example, the spectral subspace of $\Delta_{\mathbb{L}}$ with energy $E_{F}$ (“flat band eigenspace”) contains compactly supported states, and any flat band eigenstate can be approximated by such compactly supported states [Kuchment2005].
Among these compactly supported states is a family of homogeneous density, alternating sign, compact states. This property can be used to obtain lattices with flat-bands via a line graph construction [Mielke1991, Mielke1991a, Kollar2020].
@@
    There is a symmetry $\mathfrak{S}\colon\mathbb{L}\to\mathbb{L}$, cyclic of even order $s$, and a point $x_{0}\in\mathbb{L}$ such that ${\rm supp}(\psi)$ consists of the $s$ distinct points: $x_{0},\mathfrak{S}x_{0},\dots,\mathfrak{S}^{s-1}x_{0}$. We call this set the orbit of $x_{0}$, ${\rm Orb_{\mathfrak{S}}}(x_{0})$, under the symmetry $\mathfrak{S}$.
@@
    $\psi$ has uniform amplitude along its support, and $\psi(x)=a\sum_{j=0}^{s-1}(-1)^{j}\delta_{{\mathfrak{S}}^{j}x_{0}}(x)$ for some $a\neq 0$, and $\delta_{y}(x)$ denotes the Kronecker delta function at $y$.
@@
a rank one perturbation of $L_{{\rm bulk},a}$ on $l_{\omega^{n}}^{2}(\mathbb{L})$, given in (8). Moreover, since $\Delta_{\mathbb{L}}-E_{F}$ is negative semi-definite, each $L_{+,\omega^{n}}$ has at most one eigenvalue, denoted $\lambda_{n}(|a|^{2\sigma})$, above energy $|a|^{2\sigma}$, the upper edge of the continuous spectrum of $L_{+,\omega^{n}}$. Note that, reflecting its oscillatory nature, $\psi_{a}$ itself lies in the subspace $l_{-1}^{1}(\mathbb{L})$, i.e., $n=s/2$, and is an eigenstate of $L_{+,-1}$, with eigenvalue $2\sigma|a|^{2\sigma}$.
@@
To conclude the discussion of Theorem 1, we argue that for generic lattices there is always a $\sigma_{\mathbb{L}}^{\rm cr}>0$. In all examples known to us, and generically we expect: $0\leq\|P_{E_{F}}^{-\Delta_{\mathbb{L}}}e_{n}\|_{2}^{2}<1$, i.e., $e_{n}$ is not a flat-band eigenstate, for every $n\neq s/2$. This condition is independent of the nonlinearity parameter $\sigma$. Hence, for small enough $\sigma>0$, (11) is always guaranteed to be satisfied, therefore establishing the existence of $\sigma_{\mathbb{L}}^{\rm cr}>0$.

### v2

@@ Abstract
It is therefore particularly relevant to study the types of stable or long-lived localized excitations that can be supported in nonlinear wave models, whose underlying linear spectrum has a flat band. A prototypical example arises in the discrete nonlinear Schrödinger (DNLS) type wave systems arising in optics, superfluids and materials (see e.g. [Lederer2008_PhysRep_DiscreteSolitons, Kevrekidis2009_DNLS, malomed2013spontaneous, AblowitzCole2022_PhysicaD_Review, granularBook]), on the Kagomé lattice [Law]. Numerical studies reveal compactly supported nonlinear bound states from which one observes symmetry-breaking bifurcations  [vicencio2013discrete]. While there is substantial analytical theory of solitary wave stability, results on nonlinear dynamics of coherent structures in systems with linear flat band spectrum are mainly based on direct numerical simulations and linearized approximations [Flach:2018]. A general, unifying theory of the dynamical stability of the nonlinear states of such systems, is of potential broad interest and our goal in this Letter is to present such a theory.
@@
The band structure contains a flat band if for some $j_{0}\in\{1,\dots,N\}$, $E_{j_{0}}(k)=E_{F}$ is a constant for all $k\in\mathcal{B}$. In this paper, we focus on the setting where $\Delta_{\mathbb{L}}$ has a single flat band at the upper limit, $E_{F}$, of its spectrum. The general properties of linear operators with flat bands have been studied extensively. For example, the spectral subspace of $\Delta_{\mathbb{L}}$ with energy $E_{F}$ (“flat band eigenspace”) contains compactly supported states, and any flat band $l^{2}(\mathbb{L})$ eigenstate can be approximated by such compactly supported states [Kuchment2005]. Among these compactly supported states is a family of homogeneous density, alternating sign, compact states. This property can be used to obtain lattices with flat-bands via a line graph construction [Mielke1991, Mielke1991a, Kollar2020].
@@
    Symmetry: There is a symmetry $\mathfrak{S}\colon\mathbb{L}\to\mathbb{L}$, cyclic of even order $s$, such that $w_{\mathfrak{S}x,\mathfrak{S}y}=w_{xy}$, and a point $x_{0}\in\mathbb{L}$ such that ${\rm supp}(\psi)$ consists of the $s$ distinct points: $x_{0},\mathfrak{S}x_{0},\dots,\mathfrak{S}^{s-1}x_{0}$. ${\rm Orb_{\mathfrak{S}}}(x_{0})$ is called the orbit of $x_{0}$, under $\mathfrak{S}$.
@@
    Alternating: $\psi$ has uniform amplitude along its support, and $\psi(x)=a\sum_{j=0}^{s-1}(-1)^{j}\delta_{{\mathfrak{S}}^{j}x_{0}}(x)$ for some $a\neq 0$, and $\delta_{y}(x)$ denotes the Kronecker delta function at $y$.
  3. 3.
    Minimality: If $\Delta_{\mathbb{L}}e=E_{F}e$ and ${\rm supp}(e)\subset{\rm supp}(\psi)$, then $e\in{\rm span}\{\psi\}$.
@@
a rank one perturbation of $L_{{\rm bulk},a}$ on $l_{\omega^{n}}^{2}(\mathbb{L})$, given in (8). Moreover, since $\Delta_{\mathbb{L}}-E_{F}$ is negative semi-definite, each $L_{+,\omega^{n}}$ has at most one eigenvalue, denoted $\lambda_{n}(|a|^{2\sigma})$, above energy $|a|^{2\sigma}$, the upper edge of the continuous spectrum of $L_{+,\omega^{n}}$. Note that, reflecting its oscillatory nature, $\psi_{a}$ itself is a multiple of $e_{s/2}$, and is an eigenstate of $L_{+,-1}$, with eigenvalue $2\sigma|a|^{2\sigma}$.
@@
To conclude the discussion of Theorem 1, we argue that for generic lattices there is always a $\sigma_{\mathbb{L}}^{\rm cr}>0$. By the minimality of MCS states, $0\leq\|P_{E_{F}}^{-\Delta_{\mathbb{L}}}e_{n}\|_{2}^{2}<1$, i.e., $e_{n}$ is not a flat-band eigenstate, for every $n\neq s/2$. This condition is independent of the nonlinearity parameter $\sigma$. Hence, for small enough $\sigma>0$, (11) is always guaranteed to be satisfied, therefore establishing the existence of $\sigma_{\mathbb{L}}^{\rm cr}>0$.